OpenStudy (anonymous):

Tracy ran for seven-twelfths hours and Reina ran for one-third hours. How many more hours did Tracy run? one-fourth 1 one-third one-sixth

OpenStudy (anonymous):

\[\frac{7}{12}-\frac{1}{3}=\frac{7}{12}-\frac{4}{12}\] is a start